OpenStudy (markiplier):

Which expression is equivalent to |x| < 11?

OpenStudy (kendricklamar2014):

Let's solve your inequality step-by-step. |x|<11 Solve Absolute Value. |x|<11 We know x<11 and x>−11

OpenStudy (kendricklamar2014):

x<11(Condition 1) x>−11(Condition 2) Answer: x<11 and x>−11

OpenStudy (kendricklamar2014):

yeah, and by the way this is History, u should out this in Mathematics next time (just a heads up)

OpenStudy (kendricklamar2014):

Let's solve your equation step-by-step. |x+5|=9 Solve Absolute Value. |x+5|=9 We know either x+5=9 or x+5=−9

OpenStudy (kendricklamar2014):

x+5=9(Possibility 1) x+5−5=9−5(Subtract 5 from both sides) x=4

OpenStudy (kendricklamar2014):

x+5=−9(Possibility 2) x+5−5=−9−5(Subtract 5 from both sides) x=−14

OpenStudy (kendricklamar2014):

Answer: x=4 or x=−14

OpenStudy (kendricklamar2014):

ok

OpenStudy (anonymous):

ok :)

OpenStudy (kendricklamar2014):

do u have asnwer choices??

OpenStudy (anonymous):

-5 < x < 3

OpenStudy (anonymous):

lolz

OpenStudy (kendricklamar2014):

ill give her a medal :)

OpenStudy (anonymous):

thx

OpenStudy (kendricklamar2014):

ur welcome

OpenStudy (kendricklamar2014):

1 more??

OpenStudy (anonymous):

that will take me some time but I gtg :( im so sorry!